Starting from rest, a truck travels in a straight line for 9.0 s with a uniform acceleration of +1.7 m/s2. The driver then applies the brakes for 3.0 s, causing a uniform acceleration of −2.0 m/s2 over that time?

(a) What is the truck's speed at the end of the braking period?
m/s

(b) What is the total distance traveled by the truck (from the point where it started at rest to the end of the braking period)?
in meters
Answer #1

a)

Consider the motion of truck while it accelerate :

vo = initial velocity = 0 m/s

a = acceleration = 1.7 m/s2

t = time period = 9 sec

v = final velocity after 9 sec .

using the kinematics equation

v = vo + a t

v = 0 + (1.7) (9)

v = 15.3 m/s

Consider the motion of truck while it slows down :

vo = initial velocity = 15.3 m/s

a = acceleration = - 2 m/s2

t = time period = 3 sec

v' = final velocity after 3 sec .

using the kinematics equation

v' = vo + a t

v' = 15.3 + (- 2) (3)

v' = 9.3 m/s

b)

Consider the motion of truck while it accelerate :

vo = initial velocity = 0 m/s

a = acceleration = 1.7 m/s2

t = time period = 9 sec

d = distance traveled

using the equation

d = vo t + (0.5) a t2

d = (0) (9) + (0.5) (1.7) (9)2

d = 68.85 m

Consider the motion of truck while it slows down :

vo = initial velocity = 15.3 m/s

a = acceleration = - 2 m/s2

t = time period = 3 sec

d' = distance traveled

using the equation

d' = vo t + (0.5) a t2

d' = (15.3) (3) + (0.5) (- 2 ) (3)2

d' = 36.9 m

D = Total distance = d + d = 68.85 + 36.9 = 105.75 m